OpenStudy (anonymous):

A sign above a sweater says “Take an additional 50% off.” The price tag shows a price that is 40% of the regular price. What percent of the regular price is the discounted price (the price after both discounts)?

OpenStudy (anonymous):

I think the answer might be 30% I just don't know how to get that answer.

OpenStudy (anonymous):

the price is x you would take 40% off then another 50% so total u would take off 90%

OpenStudy (anonymous):

so that would leave you 10%

myininaya (myininaya):

lets think about it. Say we have something that cost 5 dollars regular price. If we take 40% off, we get the discount of 2 dollars. We are now paying 3 dollars for the item. But there is another discount of 50%. So the discounted price is is a $1.50. 1.50/5.00=.3=30% is what we are paying. The discount is 70%.

OpenStudy (anonymous):

The price would be 20% of the regular price. Let P = regular price, R = rack price, F = final price. \[ R = \frac{40}{100}*P\] \[ F = \frac{50}{100}*R \rightarrow F= \frac{50}{100}*\frac{40}{100} * P \] \[= \frac{2000}{10000}*P = \frac{20}{100}*P = \text{20% of the regular price}\]

myininaya (myininaya):

no its 30% is what we are paying and 70% is the discount I'm 100% sure

myininaya (myininaya):

x represents the price of an item before any discounts. the price we are paying is (x-.4x)-[(x-.4x)-.5(x-.4x)]=.3x

myininaya (myininaya):

the discount price is .7x

myininaya (myininaya):

So we are paying 30%

OpenStudy (anonymous):

\[\text{"40% of the regular price"} \ne \text{"40% off the regular price"}\]

OpenStudy (anonymous):

polpak can you help me with a problem when you finish this up...

OpenStudy (anonymous):

What problem?

OpenStudy (anonymous):

A fence is to be built to enclose a rectangular area of 280 square feet. The fence along three sides is to be made of material that costs 5 dollars per foot, and the material for the fourth side costs 15 dollars per foot. Find the dimensions of the enclosure that is most economical to construct.

OpenStudy (anonymous):

Thanks so much Myininaya!! :)